OpenStudy (anonymous):

What is the solution to x=sqrt 3x plus 10? 3x+10 is all under the sqrt

Miracrown (miracrown):

\[x = \sqrt{3x + 10}\]

OpenStudy (misty1212):

HI!!

Miracrown (miracrown):

Any ideas on what to do first?

OpenStudy (anonymous):

I squared both sides and got x^2=3x+10 then took 10 from both sides and got x^2 -10=3x now I'm lost

OpenStudy (misty1212):

if it was me, i would guess that is what i would do first

OpenStudy (misty1212):

there are not that many perfect squares

OpenStudy (anonymous):

I know I need to divide the 3 but how do i get rid of the x^2

Miracrown (miracrown):

\[x ^{2} = 3x +10\]

OpenStudy (misty1212):

\[\sqrt{3x+10}\] what x will make it a perfect square ?

OpenStudy (misty1212):

not 1, 2 will but \(\sqrt{16}=4\) not \(2\) 3 doesn't work neither does 4 how about 5?

Miracrown (miracrown):

\[x ^{2} - 3x + 10 = 0\]

OpenStudy (misty1212):

got it on the fifth try

OpenStudy (anonymous):

oh ya get it all to one side

Miracrown (miracrown):

Yep

OpenStudy (misty1212):

everyone love algebra, no one wants to try experimenting why is that?

OpenStudy (anonymous):

but wouldn't be an negative 10

Miracrown (miracrown):

Do you know how to factor this? @costonla

Miracrown (miracrown):

Well in general when you have a ax^2+bx+c you factor it like this..

OpenStudy (anonymous):

(x-5)(x-3)

OpenStudy (anonymous):

(x+3)

Miracrown (miracrown):

\[(x - \space ) ( x + \space ) = 0\]

OpenStudy (anonymous):

x=5 and -3 right

Miracrown (miracrown):

Not quite because that would result in 15 not 10.

OpenStudy (anonymous):

scratch 3 and insert 2

Miracrown (miracrown):

The two last numbers multiplied together need to =10

Miracrown (miracrown):

\[(x-5) \space (x+2) \space = 0\]

OpenStudy (anonymous):

if i looked at the question right when we moved the 3x +10 to the other side we should have put a negative in front of both

Miracrown (miracrown):

Yeah sorry I mistyped that but fixed it

OpenStudy (anonymous):

so then x= 5 and -2

Miracrown (miracrown):

Correct!

OpenStudy (anonymous):

ok thank you, I forgot to move all to one side and set to 0

Miracrown (miracrown):

Nw :)